OpenStudy (anonymous):

WILL GIVE MEDAL Given two points, write the equation of the line in slope-intercept form (4,7),(-2,1) PLEASE SHOW WORK

OpenStudy (anonymous):

@raffle_snaffle can you help..?

OpenStudy (anonymous):

you need the slope first \[\frac{y_2-y_1}{x_2-x_1}\] which in your case is \[\frac{7-1}{4-(-2)}\] what do you get for this number?

OpenStudy (anonymous):

6/6..?

OpenStudy (anonymous):

yes, and what is \(\frac{6}{6}=?\)

OpenStudy (anonymous):

1? hahahah

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

now use the point slope formula \[y-y_1=m(x-x_1)\] with \[x_1=4,y_1=7\] and \(m=1\) so you don't really have to write \(m\) at all you get \[y-7=x-4\]

OpenStudy (anonymous):

solve for \(y\) by adding \(7\) to both sides

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

@satellite73